OpenStudy (anonymous):

An endangered species of fish has a population that is decreasing exponentially (A = ae kt). The population 9 years ago was 1700. Today, only 1000 of the fish are alive. Once the population drops below 100, the situation will be irreversible. When will this happen, according to the model? (Round to the nearest whole year.)

OpenStudy (anonymous):

So we start with:\[ A=ae^{kt} \]

OpenStudy (anonymous):

A=1000e^ (i get stuck here)

OpenStudy (anonymous):

Okay so you have: \[ A = 1000e^{kt} \]

OpenStudy (anonymous):

That was using the fact `Today, only 1000 of the fish are alive`

OpenStudy (anonymous):

Next you want to use the fact: `The population 9 years ago was 1700.`

OpenStudy (anonymous):

drops by a factor of \(\frac{10}{17}\) every 9 years if you don't want to go through all that mess, use \[\large 1000\times \left(\frac{10}{17}\right)^{\frac{t}{9}}\]

OpenStudy (anonymous):

This is saying \(t=-9 , A=1700\)

OpenStudy (anonymous):

So \[ 1700=1000e^{k(-9)} \]

OpenStudy (anonymous):

We can solve for \(k\) now: \[ 1.7 = e^{-9k}\implies \ln(1.7) =-9k \implies k = -\frac{\ln(1.7)}{9} \]

OpenStudy (anonymous):

Now we get: \[ A = 1000e^{-\ln(1.7)t/9} \]

OpenStudy (anonymous):

to get rid of the \(\ln\) we can use that: \[ e^{\ln(1.7)} = 1.7 \]This leads us to: \[ A = 1000 (1.7)^{-t/9} \]

OpenStudy (anonymous):

how do I get rid of the t? and thanks so much, this is very helpful.

OpenStudy (anonymous):

We don't want to get rid of \(t\). We want to solve for it.

OpenStudy (anonymous):

We use the final question to do so: `Once the population drops below 100, the situation will be irreversible. When will this happen, according to the model`

OpenStudy (anonymous):

This means \(A<100\).

OpenStudy (anonymous):

So we want: \[ 1000(1.7)^{-t/9} < 100 \]We need to solve it. Our answer should be along the lines of:\[ t> \ldots \]Because we know that it's from some point in time and all times beyond that into the future.

OpenStudy (anonymous):

What Im not getting is the t/9. i do not know what to do with it.

OpenStudy (anonymous):

im sorry I am asking so much questions, I am not that good at math.

OpenStudy (anonymous):

Okay, first think we do is divide simplify a bit: \[ (1.7)^{-t/9} < 0.1 \]

OpenStudy (anonymous):

Now we need to remember that we have: \[ (1.7)^{-t/9} = (1.7)^{t(-1/9)} \]

OpenStudy (anonymous):

We can get rid of this \(-1/9\) by using the exponent laws: \[ (x^{-1/9})^b = x^{-1/9 \times b} \]

OpenStudy (anonymous):

We want \[ -1/9\times b = 1 \]So solving for \(b\) gives us \(b=-9\).

OpenStudy (anonymous):

So we take both sides to the power of \(-9\): \[ [(1.7)^{t(-1/9)}]^{-9} < (0.1)^{-9} \]This simplifies to: \[ (1.7)^t < (0.1)^{-9} \]

OpenStudy (anonymous):

This next part is a bit tricky, but we have to use logarithms.

OpenStudy (anonymous):

Hold on, when I raised to the power of \(-9\), the inequality should have flipped.

OpenStudy (anonymous):

\[ (1.7)^t > (0.1)^{-9} \]Raising to negative powers flips inequalities.

OpenStudy (anonymous):

\[ \ln[(1.7)^t] > \ln[(0.1)^{-9}] \implies t\ln(1.7) > -9\ln(0.1)\implies t> -9\frac{\ln(0.1)}{\ln(1.7)} \]

OpenStudy (anonymous):

i got 39 years. I think thats wrong.

OpenStudy (anonymous):

Their solution is saying \(t>39\)

OpenStudy (anonymous):

with your help, I was right with the 39 yearss?

OpenStudy (anonymous):

I believe so.

OpenStudy (anonymous):

thanks alot:-)